OpenStudy (anonymous):

log(little 4)(x^2+5)=4

OpenStudy (anonymous):

4^4= x^2 +5 251= x^2 x=+- sqrt 251

OpenStudy (anonymous):

same as before, 4^4=x^2+5, isolate for x you can do the rest

OpenStudy (anonymous):

ahhh i m having a really bad day...making too many mistkaes...time to off to bed..good night

OpenStudy (anonymous):

thank you, i didnt know if 251 was the end

OpenStudy (anonymous):

no problem..good luck with ur maths...